A constant force of 2.50 N accelerates a stationary particle of mass 15 g through a displacement of 2.50 m.Find the work done and the average power delivered?

Grade:11

2 Answers

Amitesh Sahu
6 years ago
Work=Force×Displacement=2.5×.2.5=625JAlso,Force=Mass×AccelerationOr,2.5=0.015×a ,Or a=500/3Displacement=ut+1/2at2,Or 2.5=1/2×500/3×t2So, t=0.00173Now, Power=Work/TimeOr, P=6.25/0.0173=31.6J
Pratham
5 years ago
Work done=force*displacement for straight line motion
Work done=2.5*2.5=6.25 J
acceleration=force÷mass
Acceleration=2.5*1000÷5=500÷3
s=ut+1÷2(at^2)
2.5=0+1÷2(500÷3)t^2
>t=0.173s
 
Power=(work done)÷time=6.25÷0.173=36.1W